II. Derivations (You must show all your work for full credit.)i. Given the model y=XB+ɛ, derive the least squares estimate for ß? (10 points) ii. Show that B=(x+x)xy is an unbiased esti
Answer #1

Let

X be a nXk matrix

Y,E and \beta be nX1 vectors

X' is the transpose of X

\beta ^ is the estimate for \beta

Now,

1596222182318_image.png

or Y= X\beta+E

i. Using principle of ordinary least squares we minimise the sum of squared residuals :

E=Y- X\beta (Residuals)

Sum of squared residuals: E'E=(Y- X\beta^)' (Y- X\beta^)= Y'Y-\beta^'X'Y-Y'X\beta^+\beta^'X'X\beta^ = Y'Y-2\beta^'X'Y+\beta^'X'X\beta^

To minimise, we partially take the derivative with \beta ^

1596222561402_image.png

Now,we get the normal equation as

(X'X)\beta^=X'Y

or \beta ^=((X'X)^-1)*X'Y ---1

ii.

Taking expectations on both sides of equation 1

E( \beta ^)=E(((X'X)^-1)*X'Y)

=((X'X)^-1)X' E(Y)= ((X'X)^-1)X' (X\beta+E)= ((X'X)^-1)X'X\beta= \beta (E(E)=0)

Hence E( \beta ^)= \beta

Therefore unbiasedness is proved

iii. Variance \beta ^ = E(( \beta ^-\beta)( \beta ^-\beta)')= E( ((X'X)^-1)X'u)((X'X)^-1)X'u)')= E((X'X)^-1)X'uu'X(X'X)^-1)= (X'X)^-1)X'E(uu')X(X'X)^-1=(X'X)^-1)X'(sigma sq)X(X'X)^-1=(sigma sq)(X'X)^-1)X'X(X'X)^-1= (sigma sq)(X'X)^-1 which is the desired variance covariance matrix.